OpenStudy (anonymous):

find the inverse of f(x)=1+8x/7-3x

OpenStudy (anonymous):

Depending on the way you wrote it, it's either f^(-1) (x) = -7/13 (x-1) or f^(-1) (x) = (7 x-1)/(3 x+8)